Section 07  Lie Groups and Lie Algebras

Section 07: Lie Groups and Lie Algebras

Invited Lectures

Wednesday, 19 August  MA 005

14:00 – 14:45 Minoru Wakimoto, Kyushu University, Fukuoka, Japan Representation theory of affine superalgebras at the critical level (I: p.96) Chair: Richard E. Borcherds

15:00 – 15:45 Kari Vilonen, Brandeis University, Waltham, USA Topological Methods in Representation Theory (I: p.95) Chair: Wolfgang Soergel

Wednesday, 19 August  MA 004

16:15 – 17:00 Laurent Lafforgue, Université Paris-Sud, Orsay, France Chtoucas de Drinfeld et applications (I: p.92) Chair: James Arthur

17:15 – 18:00 Robert Edward Kottwitz, University of Chicago, Chicago, USA Harmonic analysis on semisimple p-adic Lie algebras (I: p.91) Chair: Michael Rapoport

Thursday, 20 August  MA 004

14:00 – 14:45 Alexander Eskin, University of Chicago, Chicago, USA Counting problems and semisimple groups (I: p.91) Chair: Peter Sarnak

15:00 – 15:45 Joseph Bernstein, Tel Aviv University, Tel Aviv, Israel Analytic structures on representation spaces of reductive groups (I: p.88) Chair: Wolfgang Soergel

Friday, 21 August  MA 004

16:15 – 17:00 Vera Serganova, University of California, Berkeley, USA On representations of Lie superalgebras (I: p.95) Chair: Olivier Mathieu

17:15 – 18:00 Shahar Mozes, Hebrew University, Jerusalem, Israel Products of trees, lattices and simple groups (I: p.93) Chair: Peter Sarnak

Saturday, 22 August  MA 004

14:00 – 14:45 James Greig Arthur, University of Toronto, Toronto, Canada Towards a stable trace formula (I: p.87) Chair: Michael Rapoport

15:00 – 15:45 Ivan V. Cherednik, The University of North Carolina at Chapel Hill, Chapel Hill, USA From double Hecke algebra to analysis (I: p.89) Chair: Olivier Mathieu

Lie Groups and Lie Algebras  Section 07

Short Communications and Poster Sessions

Short Communications  Thursday, 20 August
Symmetric Spaces  H 112

Chair: Walter J. Michaelis

16:15 – 16:30 Sergei S. Platonov, Petrozavodsk State University, Petrozavodsk, Russia Invariant subspaces of some function spaces on symmetric spaces (p.109)

16:30 – 16:45 Shigeru Aoki, Takushoku University, Hachioji-shi, Tokyo, Japan Invariant Eigendistributions on U(1, 2)/(U(1, 1) × U(1, 1)) (p.101)

16:45 – 17:00 Adam Koranyi, Lehman College, Tenafly, NY, USA Equivariant differential operators on homogeneous vector bundles over G/MAN (p.107)

Lie Groups  H 112

Chair: Adam Koranyi

17:15 – 17:30 Dietrich Burde, University of Düsseldorf, Düsseldorf, Germany Affine and projective structures on Lie groups (p.103)

17:30 – 17:45 Michael Wüstner, Technische Universität Darmstadt, Darmstadt, Germany Lie groups with surjective exponential function (p.111)

17:45 – 18:00 Walter J. Michaelis, University of New Orleans, LA, USA The exponential matrices of non-commuting matrices may commute (p.108)

Short Communications  Friday, 21 August
Quantum Groups and Vertex Operator Algebras  H 112

Chair: Gerald Höhn

14:00 – 14:15 Kailash C. Misra, North Carolina State University, Raleigh, NC, USA Demazure Crystals for quantum affine Lie algebras (p.108)

14:15 – 14:30 Naihuan Jing, North Carolina State University, Raleigh, NC, USA Quantized Kac-Moody algebras and symmetric functions (p.106)

14:30 – 14:45 Gerald Höhn, Universität Freiburg, Freiburg, Germany Vertex Operator Algebras and 3-Dimensional Topological Quantum Field Theories (p.105)

Quantum Groups and Poisson Homogeneous Spaces  H 112

Chair: Kailash C. Misra

15:00 – 15:15 Pu Zhang, Univ. of Sci. and Tech. of China, Hefei, PR China Ringel-Hall algebras and quantum groups (p.111)

15:15 – 15:30 Ho Hai Phung, International Centre for Theoretical Physics, Trieste, Italy Rational representations of matrix quantum groups of type A_{n} (p.109)

15:30 – 15:45 Eugene A. Karolinsky, Kharkov State University, Kharkov, Ukraine A classification of Poisson homogeneous spaces of complex reductive Poisson-Lie groups (p.106)

Short Communications  Saturday, 22 August
Topological Groups  H 112

Chair: Wolfgang Reinhold Bergmann

16:15 – 16:30 Boris S. Khots, Compressor Controls Corporation, Des Moines, Iowa, USA Fifth Hilbert problem and infinite-dimensional P-groups (p.106)

16:30 – 16:45 Grigori Lazarevich Litvinov, International Sophus Lie Center, Moscow, Russian Federation Group algebras and nonunitary representations of smooth groups (p.107)

16:45 – 17:00 Julio Cesar Rebelo, Pontifical Catholic University of Rio de Janeiro, Rio de Janeiro, Brazil Associating vector fields to certain subgroups of Diff^{w}(S^{1}): applications to ergodicity and rigidity

(p.109)

Section 07  Lie Groups and Lie Algebras

Amenable Groups  H 112

Chair: Grigori L. Litvinov

17:15 – 17:30 Wolfgang Reinhold Bergmann, Università della Basilicata, Potenza, Italy Amenable Groups and Generalized Cuntz Algebras (p.102)

17:30 – 17:45 Alexander I. Shtern [Chtern], Moscow State University, Moscow, Russia Quasi-Symmetry and Amenability (p.110)

Short Communications  Monday, 24 August
Representations I  H 112

Chair: Carina Boyallian

14:00 – 14:15 Jose I. Liberati, MIT, Cambridge, USA Quasifinite Highest Weight Modules over the Lie Algebra of Matrix Differential Operators on the Circle (p.107)

14:15 – 14:30 María L. Barberis, FaMAF — Universidad Nacional, Córdoba, Argentina Quaternionic representations (p.101)

14:30 – 14:45 Paolo Casati, University of Milan, Milan, Italy α–stratified modules of real reductive groups (p.104)

Representations II  H 112

Chair: Jose I. Liberati

15:00 – 15:15 Leandro R. Cagliero, Fa.M.A.F. — Univ. Nac. de Cordoba, Cordoba, Argentina (K, M)-spherical submodules of the universal enveloping algebra of a real rank one semisimple Lie group G_{∘} (p.103)

15:15 – 15:30 Carina Boyallian, MIT-Math. Dept., Cambridge, MA, USA On Bernstein polynomials for compact Lie groups (p.103)

15:30 – 15:45 Tuong Ton-That, University of Iowa, Iowa City, USA I Invariant Theory for some Infinite-Dimensional Lie Groups (p.110)

Short Communications  Tuesday, 25 August
Various Topics I  H 112

Chair: Michael Forger

16:15 – 16:30 Lizhen Ji, University of Michigan, Ann Arbor, MI, USA The trace class conjecture in the Selberg trace formula (p.106)

16:30 – 16:45 Alexander Baranov, Inst. Math. Acad. Sci. Belarus, Minsk, Belarus Finitary simple Lie algebras (p.101)

16:45 – 17:00 Michael Forger, Universidade de São Paulo, São Paulo, Brazil Symmetry and Symmetry Breaking: An Algebraic Approach to the Genetic Code (p.104)

Various Topics II  H 112

Chair: Alexander Baranov

17:15 – 17:30 Thomas Bouetou, Université de Yaounde I, Yaounde, Cameroun About Bol algebras (p.102)

17:30 – 17:45 Lyudmyla L. Barannyk, National Academy of Sciences of Ukraine, Kyiv, Ukraine On an algorithm for classifying subalgebras of the Poincaré algebra AP (2, n)

Poster Session  Monday, 24 August 16:15 – 18:00  MA 144

Helge Glöckner, Technische Universität Darmstadt, Darmstadt, Germany Scale functions on p-adic Lie groups (p.105)

Nikolai Gromov, Komi Science Center RAS, Syktyvkar, Russia Nonsemisimple Cayley-Klein groups and its quantization (p.105)

Lie Groups and Lie Algebras  Section 07

Duncan J. Melville, St. Lawrence University, Canton, NY, USA

A survey of quantum deformation results (p.108) Victor M. Petrogradsky, Ulyanovsk State University, Ulyanovsk, Russia

On subalgebras of free Lie algebras (p.108) Jorge Soto-Andrade, University of Chile, Santiago, Chile

Gel’fand Models, Weil Representations, and Twisted Spherical Functions (p.110)